# Less is More: Skim Transformer for Light Field Image Super-resolution

### v2

@@ III-C Skim Transformer
$\mathcal{Q}_{i}=Q_{i}(\mathcal{D}_{i}),\;\mathcal{K}_{i}=K_{i}(\mathcal{D}_{i}),\;\mathcal{V}_{i}=X_{i},\;$
$\mathcal{D}_{i}=D_{i}(\bar{X}_{i});$
$X_{i}\in\mathbb{R}^{U\times V\times W\times H\times C/N_{DSA}};\;$
$\quad\bar{X}_{i}\in\mathbb{R}^{WH\times\mathcal{S}_{i}C/N_{DSA}},\;\{\mathcal{S}_{i}\}\subseteq\{U\times V\};$
$Q_{i},K_{i}:\mathbb{R}^{C_{D}}\mapsto\mathbb{R}^{C_{QK}},D_{i}:\mathbb{R}^{\mathcal{S}_{i}C/N_{DSA}}\mapsto\mathbb{R}^{C_{D}}.$
With the obtained query, key, and value matrices, the upcoming self-attention calculation, including attention score matrix and self-attention output, follows the typical Transformer method:
$DSA_{i}(\mathcal{Q}_{i},\mathcal{K}_{i},\mathcal{V}_{i})=Softmax(\mathcal{Q}_{i}\mathcal{K}_{i}^{T}/\sqrt{C_{QK}})\mathcal{V}_{i},$
where $DSA_{i}\in\mathbb{R}^{WH\times UVC/N_{DSA}}$. Note that as the channel number is made consistent as $C/N_{DSA}$ through the Skim Transformer, the shape of self-attention output aligns with the input, therefore, channel projection is not necessary, and the feed-forward network can be removed to reduce the computation.
@@
$\mathbf{Y}=SkimT(\mathbf{X})=[DSA_{1},DSA_{2},\cdots,DSA_{N_{DSA}}]$
where [][$\cdot$][ ⋅ ] signifies the concatenation operation, and $\mathbf{Y}\in\mathbb{R}^{WH\times UVC}$ and finally $\mathbb{R}^{U\times V\times W\times H\times C}$.
@@ III-D Discussion
With these innovations, the Skim Transformer confers three principal advantages:
@@
First, Skim Transformers systematically disentangle the process of disparity modeling via a divide-and-conquer strategy. It splits the task into multiple DSA branches, each serving as a specialized prior and constructing its own attention score matrix over a selectively skimmed SAI subset in the query and key matrices to target specific disparity ranges. By ensuring each branch operates separately, any specific disparity, especially those underrepresented in training samples yet critical for high-fidelity reconstructions, is preserved from being overlooked and suppressed. This architecture enables SkimLFSR to become inherently disparity-aware. Further analysis in Sections IV-E1 and IV-E2 investigates how this design correlates with scene depth and camera configurations in LF images. Although these two modalities are not explicitly provided during regression-based training, SkimLFSR appears to develop a discriminative tendency aligned with them.
@@
Second, the Skim Transformer offers flexibility in the construction of skimmed SAI sets, enabling customized attention for any desired disparity range. As illustrated in Fig.3, the $DSA_{1}$ branch leverages the outer SAIs at the corners to model larger disparities, while $DSA_{2}$ employs inner SAIs for shorter disparity modeling. Both the number and spatial configuration of SAIs can vary between branches and may overlap. Regardless of the configuration, the resulting tokens remain in the spatial domain $(W\times H)$, ensuring that the computed attention matrices maintain a shape of $(WH\times WH)$ before being applied to the value matrix. This flexibility allows the network for meeting diverse requirements and scenarios. In Section IV-E3, we experiment with a variety of SAI combinations to showcase this flexibility and analyze the rationale behind SAI selection in maximizing the performance.
$\mathcal{C}(T)$ $\mathcal{C}(SkimT)$ $\mathcal{C}(SkimT)/\mathcal{C}(T)$ [Symbolic] $\mathcal{C}(SkimT)/\mathcal{C}(T)$ [Numeric]
[…]
    Note: $\mathcal{C}(T)$ and $\mathcal{C}(SkimT)$ denotes the computation complexity of a conventional spatial Transformer and a Skim Transformer, respectively.
@@
Third, the Skim Transformer significantly reduces computational redundancy without information loss. This is achieved by applying the skimming process only to the query $\mathcal{Q}_{i}$ and key $\mathcal{K}_{i}$ matrices to compute the attention score matrix, while the value matrix $\mathcal{V}_{i}$ retains the full set of SAIs. As a result, the revised self-attention mechanism still preserves full access to LF information while substantially lowering computational cost. In other words, the principle of “less” is realized not at the expense of losing information.
@@
Table I presents a theoretical computation complexity comparison between a Skim Transformer ($SkimT$) and a conventional spatial Transformer ($T$) [25]. The overall complexity is broken down into four components: matrix projection, attention score matrix, self-attention output, and feed-forward network.
Assuming $U=V=5$, $\mathcal{S}_{i}=4$, $C_{D}=2C$, $C_{QK}=C$, and $N_{DSA}=2$ (as used in the experiments), the Skim Transformer substantially brings computation complexity down as follows:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
    Self-attention Output: The self-attention output involves a matrix multiplication between the attention score matrix and the value matrix. This step is identical for both the spatial Transformer and the Skim Transformer, and thus they share the same computational complexity $\mathcal{O}(UV(WH)^{2}C)$ in this part.
  4. 4.
    Feed-forward Network: Since the shape of the Skim Transformer’s self-attention output matches its input, there is no need for a feed-forward network for channel projection, unlike in conventional spatial Transformers. This leads to a complete computation reduction from $\mathcal{O}(UVWHC^{2})$ to $\mathcal{O}(0)$.
These optimizations contribute to the Skim Transformer’s substantial improvement in efficiency, as further demonstrated in Section IV-D and Fig. 6.
IV Experiments
IV-A Experimental Settings
We conduct experiments using the widely adopted BasicLFSR framework [67] and following its protocol. It consists of five datasets: EPFL [68], HCInew [69], HCIold [70], INRIA [71], and STFgantry [21]. These datasets are split into 70/20/10/35/9 training samples and 10/4/2/5/2 testing samples, respectively. As per the protocol [67], only the central $\times 5$ 5 SAIs are used. During training, each SAI is divided into patches of size $\times 64$ 64 or $\times 128$ 128 to form HR patches, and the corresponding LR patches are generated via bicubic down-sampling by factors of 1/21/21 / 2 or 1/41/41 / 4, corresponding to the 2×2$\times$2 × and 4×4$\times$4 × LFSR scales. We adopt the $L_{1}$ loss function and the Adam optimizer with a learning rate of $2\times 10^{-4}$ and a batch size of 4. The network is trained for 100 epochs, followed by an additional 10 epochs for fine-tuning with a learning rate of $2\times 10^{-5}$.
Regarding SkimLFSR’s hyperparameters, we empirically set the feature channels $C=48$ across all Transformers and convolutions, and $C_{D}=96$ and $C_{QK}=48$ for Skim Transformers. Within each Skim Transformer, as shown in Fig. 3, we employ two DSA branches ($N_{DSA}=2$), one of which leverages the four outer SAIs (blue) representing the largest disparity range, and the other one uses the four inner SAIs (red) closer to the center for the smaller disparity range. Due to the improved efficiency of Skim Transformers, we can construct a deeper network with 20 correlation blocks ($N_{CB}=20$) for both the 2×2$\times$2 × and 4×4$\times$4 × LFSR scale.
IV-B Quantitative Comparison
@@
Methods
EPFL
(70/10)
HCInew
(20/4)
HCIold
(10/2)
INRIA
(35/5)
STFgantry
(9/2)
Average
@@ Methods
+0.21/+0.0024 +0.27/+0.0025 +0.28/+0.0013 +0.15/+0.0009 +0.80/+0.0046 +0.35/+0.0023
    Note: The number pair below each dataset represents the number of training/testing samples.
A quantitative comparison is conducted to compare the overall performance of SkimLFSR with state-of-the-art methods at the 2×2$\times$2 × and 4×4$\times$4 × scales. PSNR and SSIM are calculated on the Y channel of the YCbCr color space as the quantitative metrics of model performance.
The comparative evaluation encompasses both convolution-based and Transformer-based approaches. The convolution-based methods include LFSSR [53], LF-ATO [54], LF-InterNet [55], LF-IINet [56], DKNet [59], LFSSR-SAV [58], DistgSSR [60], and HLFSR [61]. The Transformer-based methods comprise DPT [62], LFT [26], EPIT [24], LF-DET [25], and M2MT-Net [27].
The quantitative results presented in Table II demonstrate that SkimLFSR achieves superior performance across both 2×2$\times$2 × and 4×4$\times$4 × scales on all five datasets. Specifically, SkimLFSR attains an average performance advantage of 0.59 dB and 0.35 dB PSNR over the second-best M2MT-Net at 2×2$\times$2 × and 4×4$\times$4 × scales, respectively. When examining individual datasets, the performance gains range from a minimum of 0.28 dB PSNR on the HCIold dataset at the 2×2$\times$2 × scale and 0.15 dB PSNR on INRIA at the 4×4$\times$4 × scale. A particularly strong performance is observed at the 2×2$\times$2 × scale on the EPFL dataset, with SkimLFSR achieving a gain of 0.44 dB, and an even greater improvement of 0.80 dB on the STFgantry dataset at the 4×4$\times$4 × scale. This substantial improvement aligns with our hypothesis that the disparity disentanglement capability of Skim Transformers enables SkimLFSR to effectively handle diverse disparity ranges and deliver a balanced performance. This advantage is particularly pronounced on the severely underrepresented STFgantry dataset as it features larger disparity ranges but contains only nine training samples, representing merely 6.25% of the total 144 training samples, making it the most data-scarce subset in our evaluation.
(a) 2×2$\times$2 × LFSR (b) 4×4$\times$4 × LFSR
@@
To provide additional insight into SkimLFSR’s comparative advantages, we present a radar plot in Fig. 4. At the 2×2$\times$2 × scale, although EPIT exhibits strong performance on the large-disparity STFgantry dataset, this appears to result from overfitting to large disparity ranges, as evidenced by its substantial performance degradation on other datasets featuring smaller disparity ranges, such as EPFL and INRIA. Similarly, at the 4×4$\times$4 × scale, while M2MT-Net demonstrates competitive performance across most datasets, it achieves only limited gains on STFgantry. These observations highlight the challenge faced by existing LFSR networks in maintaining consistently strong performance across various disparity ranges. In contrast, no competing method matches our SkimLFSR’s performance on any individual dataset, underscoring its robustness and ability to effectively manage the balance of diverse disparity conditions.
IV-C Qualitative Comparison
@@
Figure 5: Qualitative comparison of 4×4$\times$4 × LFSR on samples. For each sample, the first row displays the ground-truth image and corresponding error maps for each method. The second row presents zoom-in views of regions highlighted by blue and red bounding boxes, accompanied by their respective EPIs. The third row shows error maps of the zoom-in regions. The bottom row shows the sample name and PSNR/SSIM values.
@@
We showcase SkimLFSR’s superior performance through qualitative evaluations at the 4×4$\times$4 × scale in Fig. 5.
@@
The first example, shown in (a) Perforated_Metal_3, is a real LF image from the EPFL dataset captured with a Lytro Illum camera. This challenging scene features complex occlusion patterns with dense perforations and bright light spots. SkimLFSR’s reconstruction exhibits remarkable detail preservation, as evidenced in the blue region where light spot details emerge clearly, and in the red region where both light spots and hole edges are sharply delineated. In contrast, competing methods suffer from light leakage artifacts around bright spots and produce indistinct hole boundaries. The corresponding EPIs further confirm that SkimLFSR maintains superior detail fidelity and angular consistency compared to alternative approaches. As a result, SkimLFSR achieves a 0.40 dB PSNR improvement over the second-best M2MT-Net.
@@
Similar improvements are observed in (b) Building_Decoded from the INRIA dataset, also captured with a Lytro Illum camera. SkimLFSR attains a 0.23 dB PSNR gain over the runner-up M2MT-Net. The error maps within the highlighted regions reveal that this performance enhancement stems from more accurate edge reconstruction and reduced artifacts along building window boundaries. The zoom-in views and EPIs corroborate these findings, showing notably sharper window edges in SkimLFSR’s output.
@@
These examples collectively validate SkimLFSR’s exceptional ability to reconstruct fine details across diverse disparity ranges and structural characteristics, maintaining consistent performance advantages in both foreground and background regions.
IV-D Model Efficiency
@@
TABLE III: Model efficiency and performance of SkimLFSR and state-of-the-art methods at the 4×4$\times$4 × scale. The best results are highlighted in bold.
Method
#Params.
FLOPs
Time
PSNR/SSIM
LFSSR-SAV [58] 1.54 99.45 3.20 32.19/0.9439
DistgSSR [60] 3.58 65.26 1.89 32.12/0.9439
HLFSR [61] 13.86 45.73 6.83 32.28/0.9455
LFT [26] 1.16 30.42 5.72 32.24/0.9436
EPIT [24] 1.47 71.15 2.54 32.41/0.9452
LF-DET [25] 1.68 51.20 4.90 32.41/0.9455
M2MT-Net [27] 3.99 33.85 2.13 32.71/0.9489
SkimLFSR (Ours)
$N_{CB}=4$ 0.62 17.85 1.38 32.47/0.9456
$N_{CB}=6$ 0.87 25.39 1.84 32.58/0.9474
$N_{CB}=8$ 1.13 32.94 2.31 32.80/0.9486
$N_{CB}=10$ 1.39 40.49 2.78 32.82/0.9489
$N_{CB}=12$ 1.64 48.04 3.25 32.82/0.9491
$N_{CB}=14$ 1.90 55.58 3.71 32.84/0.9499
$N_{CB}=16$ 2.16 63.13 4.18 32.91/0.9499
$N_{CB}=18$ 2.42 70.68 4.66 32.96/0.9508
$N_{CB}=20$ 2.67 78.23 5.12 33.06/0.9512

### v3

@@ III-C Skim Transformer
$\mathcal{Q}_{i}=Q_{i}(\mathcal{D}_{i}),\;\mathcal{K}_{i}=K_{i}(\mathcal{D}_{i}),\;\mathcal{V}_{i}=X_{i},\;$ (2)
$\mathcal{D}_{i}=D_{i}(\bar{X}_{i});$ (3)
$X_{i}\in\mathbb{R}^{U\times V\times W\times H\times C/N_{DSA}};\;$ (4)
$\quad\bar{X}_{i}\in\mathbb{R}^{WH\times\mathcal{S}_{i}C/N_{DSA}},\;\{\mathcal{S}_{i}\}\subseteq\{U\times V\};$ (5)
$Q_{i},K_{i}:\mathbb{R}^{C_{D}}\mapsto\mathbb{R}^{C_{QK}},$ (6)
$D_{i}:\mathbb{R}^{\mathcal{S}_{i}C/N_{DSA}}\mapsto\mathbb{R}^{C_{D}}.$ (7)
With the obtained query, key, and value matrices, the upcoming self-attention calculation, including the attention score matrix and the self-attention output, follows the typical Transformer method:
$DSA_{i}(\mathcal{Q}_{i},\mathcal{K}_{i},\mathcal{V}_{i})=Softmax(\mathcal{Q}_{i}\mathcal{K}_{i}^{T}/\sqrt{C_{QK}})\mathcal{V}_{i},$ (8)
where $DSA_{i}\in\mathbb{R}^{WH\times UVC/N_{DSA}}$. Note that as the channel number is made consistent as $C/N_{DSA}$ through the Skim Transformer, the shape of self-attention output aligns with the input. Therefore, channel projection is not necessary, and the feed-forward network can be removed to reduce the computation.
@@
$\mathbf{Y}=SkimT(\mathbf{X})=[DSA_{1},DSA_{2},\cdots,DSA_{N_{DSA}}]$ (9)
where $\cdot$] signifies the concatenation operation, and $\mathbf{Y}\in\mathbb{R}^{WH\times UVC}$ and finally $\mathbb{R}^{U\times V\times W\times H\times C}$.
$\mathcal{O}(T)$ $\mathcal{O}(SkimT)$ $\mathcal{O}(SkimT)/\mathcal{O}(T)$ [Symbolic] $\mathcal{O}(SkimT)/\mathcal{O}(T)$ [Numeric]
(1) Matrix Projection $3UVWHC^{2}$ $\sum_{i=1}^{N_{DSA}}\mathcal{S}_{i}WH(\frac{C}{N_{DSA}}C_{D}+2C_{D}C_{QK})$ $\frac{\sum_{i=1}^{N_{DSA}}\mathcal{S}_{i}(C_{D}C/N_{DSA}+2C_{D}C_{QK})}{3UVC^{2}}$ $40:75\approx 53\%$
(2) Attention Score Matrix $UV(WH)^{2}C$ $N_{DSA}(WH)^{2}C_{QK}$ $\frac{N_{DSA}C_{QK}}{UVC}$ $2:25\approx 8\%$
(3) Self-attention Output $UV(WH)^{2}C$ $UV(WH)^{2}C$ $1:1$ $1:1=100\%$
(4) Feed-forward Network $UVWHC^{2}$ 0 $0:1$ $0:1=0\%$
@@ III-D Discussion
With these innovations, the Skim Transformer confers four principal advantages:
III-D1 Disparity Disentanglement
@@
Skim Transformers systematically disentangle the process of disparity modeling via a divide-and-conquer strategy. It splits the task into multiple DSA branches, each serving as a specialized prior and constructing its own attention score matrix over a selectively skimmed SAI subset in the query and key matrices to target specific disparity ranges. By ensuring each branch operates separately, any specific disparity, especially those underrepresented in training samples yet critical for high-fidelity reconstructions, is preserved from being overlooked and suppressed. This architecture enables SkimLFSR to become inherently disparity-aware. Further analysis in Sections IV-E1 and IV-E2 investigates how this design correlates with scene depth and camera configurations in LF images. Although these two modalities are not explicitly provided during regression-based training, SkimLFSR appears to develop a discriminative tendency aligned with them.
III-D2 Flexibility in Skimmed SAI Sets
@@
Skim Transformers offer flexibility in the construction of skimmed SAI sets, enabling customized attention for any desired disparity range. As illustrated in Fig.3, the $DSA_{1}$ branch leverages the outer SAIs at the corners to model larger disparities, while $DSA_{2}$ employs inner SAIs for shorter disparity modeling. Both the number and spatial configuration of SAIs can vary between branches and may overlap. Regardless of the configuration, the resulting tokens remain in the spatial domain $(W\times H)$, ensuring that the computed attention matrices maintain a shape of $(WH\times WH)$ before being applied to the value matrix. This flexibility allows the network to meet diverse requirements and scenarios. In Section IV-E3, we experiment with a variety of SAI combinations to showcase this flexibility and analyze the rationale behind SAI selection in maximizing the performance.
III-D3 Lower Computational Complexity
@@
Skim Transformers significantly reduce computational redundancy without information loss. This is achieved by applying the skimming process only to the query $\mathcal{Q}_{i}$ and key $\mathcal{K}_{i}$ matrices to compute the attention score matrix, while the value matrix $\mathcal{V}_{i}$ retains the full set of SAIs. As a result, the revised self-attention mechanism still preserves full access to LF information while substantially lowering computational cost. In other words, the principle of “less” is realized not at the expense of losing information.
@@
Table I presents a theoretical computation complexity comparison between a Skim Transformer ($\mathcal{O}(SkimT)$) and a conventional spatial Transformer ($\mathcal{O}(T)$) [9] of four common components, namely matrix projection, attention score matrix, self-attention output, and feed-forward network. Assuming $U=V=5$, $\mathcal{S}_{i}=4$, $C_{D}=2C$, $C_{QK}=C$, and $N_{DSA}=2$ (as used in the experiments), the Skim Transformer substantially brings computation complexity down to 53% for matrix projection, 8% for attention score matrix, and 0% for feed-forward network, while maintaining the same complexity for self-attention output. A detailed explanation is provided in the supplementary material.
These improvements will be further validated in Section IV-D and Fig. 6, where the Skim Transformer is shown to achieve a better trade-off between performance and efficiency.
III-D4 Cross-angular-resolution Generalization
In the LF community, the ability to generalize across different angular resolutions has been relatively underexplored. One key reason is that some existing approaches adopt an explicit encoding of the entire angular space into a fixed-size feature space. For instance, the spatial Transformers used in M2MT-Net [21] treat the spatial locations as tokens and apply a linear layer to aggregate angular information from all SAIs, i.e., $\mathbb{R}^{U\times V}\mapsto\mathbb{R}^{C_{D}}$. Similar designs are also adopted in the convolution-based DistgSSR [46] and HLFSR [40] where their EPI feature extractors employ a convolution kernel of size related to $U$ or $V$, into feature channels $C_{D}$, and their angular feature extractors further use a convolution kernel of size $(U\times V)$ to aggregate angular information from all the SAIs. These designs inherently couple the model to a specific angular resolution, since $(U\times V)$ is fixed once a model is trained, thereby preventing operation on a different angular resolution.
@@ III-D4 Cross-angular-resolution Generalization
Methods (#Train/#Test samples) EPFL (70/10) HCInew (20/4) HCIold (10/2) INRIA (35/5) STFgantry (9/2) Average
@@
+0.21/+0.0024 +0.27/+0.0025 +0.28/+0.0013 +0.11/+0.0009 +0.80/+0.0046 +0.35/+0.0023
(a) $2\times$ LFSR (b) $4\times$ LFSR
@@
In contrast, our Skim Transformer is angular-resolution-agnostic, because the disparity embedding process operates not on the complete set of SAIs but on a skimmed subset $\{\mathcal{S}_{i}\}\subseteq\{U\times V\}$, i.e. $\mathbb{R}^{\mathcal{S}_{i}}\mapsto\mathbb{R}^{C_{D}}$ as indicated as $D_{i}$ in Equation 7. The model can naturally generalize to a new resolution $U^{\prime}\times V^{\prime}$ as long as the skimmed SAI set fits within it, i.e. $\{\mathcal{S}_{i}\}\subseteq\{U^{\prime}\times V^{\prime}\}$. Consequently, this design not only disentangles disparity information itself as described earlier but also disentangles the model from any specific angular resolution, guiding it to learn angular-resolution-agnostic disparity information and achieve a better generalization capability.
This advantage will be validated in Section IV-F3 and Table IV where the SkimLFSR model trained on $5\times 5$ SAIs is tested on $7\times 7$ SAIs without any retraining and major network modifications.
IV Experiments
IV-A Experimental Settings
We conduct experiments using the widely adopted BasicLFSR framework [14] and following its protocol. It consists of five datasets: EPFL [37], HCInew [19], HCIold [48], INRIA [25], and STFgantry [39]. These datasets are split into 70/20/10/35/9 training samples and 10/4/2/5/2 testing samples, respectively. SkimLFSR is trained for 100 epochs with a learning rate of $2\times 10^{-4}$, followed by an additional 10 epochs for fine-tuning with a learning rate of $2\times 10^{-5}$.
Regarding SkimLFSR’s hyperparameters, we empirically set the feature channels $C=48$ across all Transformers and convolutions, and $C_{D}=96$ and $C_{QK}=48$ for Skim Transformers. Within each Skim Transformer, as shown in Fig. 3, we employ two DSA branches ($N_{DSA}=2$), one of which leverages the four outer SAIs (blue) representing the largest disparity range, and the other one uses the four inner SAIs (red) closer to the center for the smaller disparity range. Due to the improved efficiency of Skim Transformers, we can construct a deeper network with 20 correlation blocks ($N_{CB}=20$) for both the $2\times$ and $4\times$ LFSR scale.
IV-B Quantitative Comparison
A quantitative comparison is conducted to compare the overall performance of SkimLFSR with state-of-the-art methods at the $2\times$ and $4\times$ scales. As per the protocol [14], only the central $5\times 5$ SAIs are used. PSNR and SSIM are calculated on the Y channel of the YCbCr color space as the quantitative metrics of model performance.
The comparative evaluation encompasses both convolution-based, Transformer-based, and Mamba-based methods. The convolution-based methods include LFSSR-SAV [8], DKNet [22], DistgSSR [46], and HLFSR [40]. The Transformer-based methods comprise DPT [42], LFT [29], EPIT [28], LF-DET [9], and M2MT-Net [21]. The Mamba-based methods include LFMamba [51].
@@ IV-B Quantitative Comparison
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Figure 5: Qualitative comparison of $4\times$ LFSR on samples. For each sample, two regions are highlighted by blue and red bounding boxes. The zoom-in views, EPI, and error maps of these two regions are presented. PSNR/SSIM results of the whole LF image are given at the bottom.
The quantitative results presented in Table II demonstrate that SkimLFSR achieves superior performance across both $2\times$ and $4\times$ scales across all datasets. Specifically, SkimLFSR attains an average performance advantage of 0.63 dB and 0.35 dB PSNR over the second-best M2MT-Net at $2\times$ and $4\times$ scales, respectively. When examining individual datasets, the performance gains range from a minimum of 0.24 dB PSNR on the HCIold dataset at the $2\times$ scale and 0.11 dB PSNR on INRIA at the $4\times$ scale. A particularly strong performance is observed on the EPFL dataset at the $2\times$ scale, with SkimLFSR achieving a gain of 0.42 dB, and an even greater improvement of 0.80 dB on the STFgantry dataset at the $4\times$ scale. This substantial improvement aligns with our hypothesis that the disparity disentanglement capability of Skim Transformers enables SkimLFSR to effectively handle diverse disparity ranges and deliver a balanced performance, and this advantage is particularly pronounced on the severely underrepresented STFgantry dataset as it features larger disparity ranges but contains only nine training samples, representing merely 6.25% (9 out of 144) of the training samples, making it the most data-scarce subset in our evaluation.
To provide additional insight into SkimLFSR’s comparative advantages, we present a radar plot in Fig. 4. At the $2\times$ scale, although EPIT exhibits strong performance on the large-disparity STFgantry dataset, this appears to result from overfitting to large disparity ranges, as evidenced by its substantial performance degradation on datasets featuring smaller disparity ranges, such as EPFL and INRIA. Mixed performance is also observed for M2MT-Net and LFMamba, which achieve strong results on EPFL and INRIA but underperform on STFgantry. In comparison, our SkimLFSR achieves a consistent performance advantage across all datasets, showing no discernible shortcomings on either large- or small-disparity datasets.
At the $4\times$ scale, recent methods yield moderate improvements over earlier approaches across most datasets. Notably, STFgantry represents a clear exception to this trend, where the best compared method, M2MT-Net, achieves only a negligible improvement of 0.02 dB PSNR over the earlier method EPIT. This stark contrast highlights the fundamental challenge faced by existing LFSR networks in maintaining consistently strong performance across varying disparity ranges. In contrast, SkimLFSR achieves a sweeping win across all datasets, with a particularly significant gain of 0.80 dB PSNR on the challenging STFgantry dataset. This comprehensive and consistent superiority underscores SkimLFSR’s exceptional robustness and effectiveness in handling diverse disparity conditions.
IV-C Qualitative Comparison
@@
We showcase SkimLFSR’s superior performance through qualitative evaluations at the $4\times$ scale in Fig. 5.
@@
The first example, shown in (a) Perforated_Metal_3, is a challenging scene that features complex occlusion patterns with dense perforations and bright light spots. SkimLFSR’s reconstruction exhibits remarkable detail preservation, as evidenced in the blue region where light spot details emerge clearly, and in the red region where both light spots and hole edges are sharply delineated. In contrast, competing methods suffer from light leakage artifacts around bright spots and produce indistinct hole boundaries. The corresponding EPIs further confirm that SkimLFSR maintains superior detail fidelity and angular consistency compared to other methods.
@@
Similar improvements are observed in (b) Building_Decoded from the INRIA dataset. The error maps within the highlighted regions reveal that this performance enhancement stems from more accurate edge reconstruction and reduced artifacts along building window boundaries. The zoom-in views and EPIs corroborate these findings, showing notably sharper window edges in SkimLFSR’s output.
@@
These examples collectively validate SkimLFSR’s exceptional ability to reconstruct fine details across diverse disparity ranges and structural characteristics, yielding consistent performance advantages in both foreground and background regions.
@@
IV-D Model Efficiency